# Industrial-scale Prediction of Cement Clinker Phases using Machine Learning

### v1

@@ Abstract
Cement production, exceeding 4.1 billion tonnes and contributing 2.4 tonnes of CO2 annually, faces critical challenges in quality control and process optimization. While traditional process models for cement manufacturing are confined to steady-state conditions with limited predictive capability for mineralogical phases, modern plants operate under dynamic conditions that demand real-time quality assessment. Here, exploiting a comprehensive two-year operational dataset from an industrial cement plant, we present a machine learning framework that accurately predicts clinker mineralogy from process data. Our model achieves unprecedented prediction accuracy for major clinker phases while requiring minimal input parameters, demonstrating robust performance under varying operating conditions. Through post-hoc explainable algorithms, we interpret the hierarchical relationships between clinker oxides and phase formation, providing insights into the functioning of an otherwise black-box model. This digital twin framework can potentially enable real-time optimization of cement production, thereby providing a route toward reducing material waste and ensuring quality while reducing the associated emissions under real plant conditions. Our approach represents a significant advancement in industrial process control, offering a scalable solution for sustainable cement manufacturing.
@@
Machine learning, Clinker phases, explainable AI,
@@ 1 Introduction
Cement is the most produced human-made material in the world [1]. Global cement production reached more than $4.1$ billion tons/year in $2023$, more than doubling from 2005 levels [2, 3], with projections indicating a further increase by 2050 [3]. Moreover, each tonne of cement releases $\sim$0.6 tonne CO2, contributing to more than 8% of global carbon emissions. This unprecedented growth [4], along with high emissions, necessitates enhanced production efficiency while maintaining stringent quality standards. The cement’s performance characteristics, particularly its 28-day compressive strength, are fundamentally determined by the clinker’s mineralogical phases—alite, belite, aluminate, and ferrite. Determining the mineralogical phases directly from the clinker plays a crucial role in determining the quality of the cement produced at the plant [5].
@@ 2.1 Cement plant data
A comprehensive two-year (01/01/2020 to 31/12/2021) operational dataset from an industrial cement plant formed the foundation for our ML framework. To the best of our knowledge, this is the first time that such a large-scale data from the cement plant has been used to develop a data-driven digital twin. The data architecture comprised three distinct databases: plant configuration parameters, namely, DB0; continuous process parameters (PP), namely, DB1; and compositional analyses, namely, DB2, including fuel characteristics, kiln feed (KF), hot meal (HM), and clinker compositions including clinker oxides (CO). Figure 1(a) illustrates the plant schematic and data collection points, with material compositions determined through XRD and XRF analyses. PP comprised on 34 features, while KF, HM, and clinker comprised of 9, 7, and 12 features, respectively. Thus, the total dataset consists of potentially 59 input features with the three clinker phases as the output. Complete details of the features in DB1 and DB2 are reported in Tables 3 and  4 in A.
@@ 3 Discussion
4 Methods
4.1 Data collection
  • 1.
  • 2.
    DB1: Process parameters including stage-wise temperature and pressure profiles, O2 content, kiln feed temperature, and calciner fuel consumption rates. Complete specifications are provided in Table 3.
  • 3.
Mineral oxide compositions in KF, HM, and clinker were quantified using X-ray fluorescence (XRF) spectroscopy. The percentage of clinker phases was determined through X-ray diffraction (XRD) analysis.
  1. 1.
    KF measurement at preheater tower inlet ($t_{0}$)
  2. 2.
  3. 3.
  4. 4.
  5. 5.
4.2 Data processing
  1. 1.
  2. 2.
    Data Consistency:
    • (a)
      Consolidated duplicate entries
    • (b)
      Removed incomplete feature sets
    • (c)
  3. 3.
The outlier threshold selection balanced data retention with statistical significance. As evidenced in Fig. 1c-e, excluded data points ($<50$ per variable) deviated significantly from the two-year compositional distributions. This filtering methodology was consistently applied across all variables to prevent spurious correlations during model training. The final curated dataset comprised 8,654 clinker measurements, representing approximately 58% of the raw data. Detailed preprocessing statistics are provided in Table 5A.
4.3 ML models
4.4 Performance Metrics
(i) Mean Absolute Percentage Error (MAPE). [32]:
$MAPE=\frac{1}{n}\sum_{i=1}^{n}\frac{|y_{p}\ (i)-y_{t}\ (i)|}{y_{t}\ (i)}$ (1)
(ii) Mean Absolute Error (MAE):
$MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{p}\ (i)-y_{t}\ (i)|$ (2)
(iii) Coefficient of Determination $(R^{2})$ [33]:
$R^{2}=1-\frac{\textnormal{RSS}}{\textnormal{TSS}}$ (3)
where,
4.5 Model Optimization Protocol
Hyperparameter optimization followed a systematic protocol [20]:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Validation: Final assessment on holdout test set.
4.6 Shapley Additive Explanations (SHAP)
where:
  • 1.
    $f$: Original model architecture
  • 2.
    $X$: Complete feature set
  • 3.
    $\lvert\ddot{z}\rvert$: Cardinality of feature combinations within non-zero feature power set
  • 4.
    $n$: Simplified input feature count
  • 5.
    $\ddot{z}\subset X^{\prime}$: Vectors with non-zero entries subset to $\ddot{X}$
  • 6.
    $f_{X}(\ddot{z})$: Model prediction for given data point
  • 7.
    $f_{X}(\ddot{z}_{\backslash k})$: Model prediction excluding feature $k$ [39]
@@ Acknowledgements
The authors declare no competing interests

### v2

@@ Abstract
Cement production, exceeding 4.1 billion tonnes and contributing 2.4 tonnes of CO2 annually, faces challenges in quality control and process optimization. Traditional process models for cement manufacturing are confined to steady-state conditions with limited predictive capability for mineralogical phases. However, modern plants operate dynamically, requiring real-time quality assessment. Here, exploiting a comprehensive two-year operational dataset from an industrial cement plant, we present a machine learning framework that accurately predicts clinker mineralogy from process data. Our model achieves unprecedented accuracy for predicting major clinker phases while requiring minimal inputs, demonstrating robust performance under varying operations. Through post-hoc explainable algorithms, we interpret the hierarchical relationships between clinker oxides and phase formation, providing insights into the functioning of an otherwise black-box model. This digital twin framework can potentially enable real-time optimization of cement production, thereby providing a route toward reducing material waste and ensuring quality while reducing the associated emissions under real-world conditions.
@@
Machine learning, clinker phases, cement manufacturing, process models.
@@ 1 Introduction
Global cement production reached more than $4.1$ billion tons/year in $2023$, more than doubling from 2005 levels [1, 2], with projections indicating a further increase by 2050 [2]. Moreover, each tonne of cement releases $\sim$0.6 tonne CO2, contributing to more than 8% of global carbon emissions. This unprecedented growth [3], along with high emissions, necessitates enhanced production efficiency while maintaining stringent quality standards. The cement’s performance characteristics, particularly its 28-day compressive strength, are fundamentally determined by the clinker’s mineralogical phases—alite, belite, aluminate, and ferrite. Determining the mineralogical phases directly from the clinker plays a crucial role in determining the quality of the cement produced at the plant [4].
@@
2 Methods
2.1 Data collection
  • 1.
  • 2.
    DB1: Process parameters including stage-wise temperature and pressure profiles, O2 content, kiln feed temperature, and calciner fuel consumption rates. Complete specifications are provided in Table 3, refappendix a data .
  • 3.
Figure 1(a) illustrates the plant schematic and data collection points for PP, KF, HM and clinker composition. Mineral oxide compositions in KF, HM, and clinker were quantified using X-ray fluorescence (XRF) spectroscopy. The percentage of clinker phases was determined through X-ray diffraction (XRD) analysis.
  1. 1.
    KF measurement at preheater tower inlet ($t_{0}$)
  2. 2.
  3. 3.
  4. 4.
  5. 5.
2.2 Data processing
  1. 1.
  2. 2.
    Data Consistency:
    • (a)
      Consolidated duplicate entries
    • (b)
      Removed incomplete feature sets
    • (c)
  3. 3.
The outlier threshold selection balanced data retention with statistical significance. As evidenced in Fig. 1 b-d, excluded data points ($<50$ per variable) deviated significantly from the two-year compositional distributions. This filtering methodology was consistently applied across all variables to prevent spurious correlations during model training. The final curated dataset comprised 8,654 clinker measurements, representing approximately 58% of the raw data. Detailed preprocessing statistics are provided in Table 5A.
2.3 ML models
2.4 Performance Metrics
(i) Mean Absolute Percentage Error (MAPE). [26]:
$MAPE=\frac{1}{n}\sum_{i=1}^{n}\frac{|y_{p}\ (i)-y_{t}\ (i)|}{y_{t}\ (i)}$ (1)
(ii) Mean Absolute Error (MAE):
$MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{p}\ (i)-y_{t}\ (i)|$ (2)
(iii) Coefficient of Determination $(R^{2})$ [27]:
$R^{2}=1-\frac{\textnormal{RSS}}{\textnormal{TSS}}$ (3)
where,
2.5 Model Optimization Protocol
Hyperparameter optimization followed a systematic protocol [19]:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Validation: Final assessment on holdout test set.
2.6 Shapley Additive Explanations (SHAP)
where, $f$: Original model architecture, $X$: Complete feature set, $\lvert\ddot{z}\rvert$: Cardinality of feature combinations within non-zero feature power set, $n$: Simplified input feature count, $\ddot{z}\subset X^{\prime}$: Vectors with non-zero entries subset to $\ddot{X}$, $f_{X}(\ddot{z})$: Model prediction for given data point, $f_{X}(\ddot{z}_{\backslash k})$: Model prediction excluding feature $k$ [33]
@@ 3.1 Cement plant data
A comprehensive two-year (01/01/2020 to 31/12/2021) operational dataset from an industrial cement plant formed the foundation for our ML framework. To the best of our knowledge, this is the first time that such a large-scale data from the cement plant has been used to develop a data-driven digital twin. The data architecture comprised three distinct databases: plant configuration parameters, namely, DB0; continuous process parameters (PP), namely, DB1; and compositional analyses, namely, DB2, including fuel characteristics, kiln feed (KF), hot meal (HM), and clinker compositions including clinker oxides (CO). The material compositions were determined through XRD and XRF analyses. PP comprised on 34 features, while KF, HM, and clinker comprised of 9, 7, and 12 features, respectively. Thus, the total dataset consists of potentially 59 input features with the three clinker phases as the output. Complete details of the features in DB1 and DB2 are reported in Tables 3 and  4 in A.
@@ 4 Conclusions
Code availability
All the source codes developed as part of this work are publicly available at the following Git repository.
@@ Acknowledgements
The authors declare no competing interests.